OpenStudy (anonymous):

The sum of two numbers is 56, and their difference is 10. What are the numbers?

OpenStudy (anonymous):

That's Hard!

OpenStudy (anonymous):

There's a WIDE Variety of numbers it can be.

undeadknight26 (undeadknight26):

divied 56 in half: 56/2 = 28 28 + 28 = 58

undeadknight26 (undeadknight26):

i forgot what to do from here oi...

OpenStudy (anonymous):

Thanks For The Medal @undeadknight26

undeadknight26 (undeadknight26):

:D it was hard lol

OpenStudy (texaschic101):

x + y = 56 --->x = -y + 56 x - y = 10 -y + 56 - y = 10 -2y + 56 = 10 -2y = 10 - 56 -2y = -46 y = -46/-2 y = 23 x + y = 56 x + 23 = 56 x = 56 - 23 x = 33 check.. x + y = 56 x - y = 10 33 + 23 = 56 33 - 23 = 10 56 = 56 (correct) 10 = 10 (correct) x = 33 and y = 23

undeadknight26 (undeadknight26):

@texaschic101 FTW!!!

OpenStudy (texaschic101):

lol :)